• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Deur/LeRose

    Log entry time 04:10:42 on July16,2003

    Entry number 103955


    keyword=Septum tuning

    Comparison of focal plan quantities and y_tg for two septum settings.
    202 looks better from the point of view of acceptance (you clearly see all the holes on the beamline side) but 198 looks better in that y and phi for the center hole are more centered on the VDC.
    Offsets are pretty small anyway at 202 (4.6 mm and 3.2 mrad) and could be due to slight misalignments and imperfect positioning of the target (upstream/downstream). We're prefer 202 as the correct setting (acceptance particularly at small angles is more important).
    Consultation with J-P underway.

    Fig 1: Input current 202 A in septum (run 2301)
    Fig 2: Input current 198 A in septum (run 2308)

    FIGURE 1

    FIGURE 2